How much heat is absorbed by 60 g of copper when its temperature is
raised from 20 °C to 80 °C?(specific heat capacity of copper = 900 J/kgK)
1
Expert's answer
2020-06-15T10:21:33-0400

We can find the required quantity of heat from the formula:


"Q = mc \Delta t,""Q = 0.06 \ kg \cdot 900 \ \dfrac{J}{kg \cdot \! ^{\circ}C} \cdot (80 \ ^{\circ}C - 20 \ ^{\circ}C) = 3240 \ J."

Answer:

"Q = 3240 \ J."